# Retrieval-Augmented Generation with Conflicting Evidence

### v1

@@ 2 Related Work
Retrieval-augmented generation (RAG) has emerged as a crucial technique for enhancing the generative capabilities of language models via integrating external knowledge retrieval instead of solely relying on parametric knowledge (Lewis et al., 2020; Karpukhin et al., 2020; Cheng et al., 2021). Variants such as REALM (Guu et al., 2020), Fusion-in-Decoder (FiD) (Izacard & Grave, 2021), and RETRO (Borgeaud et al., 2022) have further optimized retrieval methods and document encoding strategies to enhance efficiency and scalability. However, these do not fully address the critical challenge of handling conflicting information across multiple retrieved documents. Prior work, such as Chen et al. (2022); Zou et al. (2024), has shown that RAG systems often propagate misinformation if retrieved documents contain errors, or arbitrarily choose an answer or use parametric knowledge to break ties when documents provide conflicting claims. Unlike existing solutions like Self-RAG (Asai et al., 2024) that use LLM-generated critiques and Astute RAG (Wang et al., 2024a) that uses the LLM’s parametric knowledge and clustering to combine retrieved documents and filter outliers or misinformation, Madam-RAG employs multi-agent debate over several rounds, wherein each agent gets the opportunity to revise its response as well as influence other agents in each round. Moreover, in contrast to Chang et al. (2024), who use a multi-agent framework solely for scoring and filtering out noisy documents, our multi-agent approach also handles other scenarios with conflicting evidence by suppressing misinformation and presenting multiple valid answers for ambiguous queries.
@@
Several benchmarks have been proposed to evaluate RAG systems under various conditions. AmbigNQ (Min et al., 2020) and AmbigDocs (Lee et al., 2024) datasets assess the ability to handle questions with multiple valid answers, with the latter extending information about ambiguous entities across multiple documents, leading us to test on AmbigDocs. However, these datasets do not consider noise or misinformation. On the other hand, the RGB dataset (Chen et al., 2024) introduces retrieval noise, testing how well models respond to partially relevant content in the retrieved documents. HAGRID (Kamalloo et al., 2023) focuses on attribution in generative QA, while CRAG (Yang et al., 2024) presents a comprehensive benchmark covering diverse question types across five domains, varying in entity popularity and temporal sensitivity. Note that each of these datasets assumes that a given question has a single correct answer and thus they do not address scenarios with multiple conflicting yet valid answers, which our work explicitly targets. We unify these separate lines of work by introducing RamDocs that includes multiple valid answers for each query (due to ambiguity) as well as uneven numbers of supporting documents and conflicting evidence from misinformed and noisy documents.
@@
Recent studies have highlighted the challenge of knowledge conflict in LLMs, where inconsistent or conflicting information may arise from either internal parametric memory or retrieved external context. A comprehensive survey by Xu et al. (2024b) categorizes knowledge conflicts into intra-context, inter-context, and parametric conflicts, and outlines the limitations of current methods in resolving them. While prior work has made strides in addressing parametric and single-document conflicts (Gao et al., 2023; Wang et al., 2024b), recent studies highlight the difficulty of resolving disagreements across multiple contexts or retrieved documents (Chen et al., 2022; Su et al., 2024). Such inter-context conflict can stem from misinformation (Pan et al., 2023a; Zhou et al., 2023) or outdated knowledge (Kasai et al., 2023), and has been shown to impair factual accuracy (Jin et al., 2024). Our Madam-RAG approach addresses this gap by modeling each document by a separate LLM agent, using multi-agent debate to identify potential ambiguity and aggregating multiple valid answers while also suppressing and resolving conflicting evidence stemming from misinformation and noisy retrieval.
@@ 3 RamDocs: Retrieval with Ambiguity & Misinformation in Documents
Measuring the performance of RAG systems in real-world settings requires assessing each system’s ability to cope with conflicting information. Moreover, different kinds of conflict demand different solutions: ambiguity should not be treated the same way as misinformation or noise; these solutions may at times conflict with each other. While existing datasets address these factors separately, in practice they will appear together and may interact in unexpected ways. This makes measuring multiple kinds of conflict in a single dataset crucial, and motivates the need for such a representative set of retrieved documents.
@@
To rigorously evaluate the robustness of current RAG systems, we present RamDocs, a dataset designed to simulate real-world retrieval challenges, including conflicting information, noise, and misinformation. Building upon AmbigDocs (Lee et al., 2024) that focuses on evaluating ambiguity by presenting multiple valid answers for a given query, we extend it by introducing additional real-world retrieval complexities. Specifically, we randomly sample a subset of between 1 to 3 correct answers for each ambiguous query from the dataset, resulting in a total of 500 queries. This approach ensures that models must not only recognize ambiguity but also operate under realistic constraints where only partial information may be retrieved.
@@
Finally, corresponding to each query, we include documents containing misinformation and noise in order to evaluate a RAG system’s resilience against unreliable content. To incorporate misinformation, we include an additional retrieved document and replace the correct answer with a misleading or incorrect entity, ensuring the misinformation blends naturally into the document context. This is similar to the approach in Longpre et al. (2021), who replace the original answer in the context with a similar type of answer, but here, we adapt the procedure to modify the retrieved documents. We randomly determine the number of misinformation documents added to each query ranging from 0-2, reflecting the unpredictable nature of misinformation in real-world retrieval. Additionally, we also introduce noisy documents for each query (sampled from [0, 2]), which are either irrelevant to the query or contain low-quality content, to assess the system’s ability to filter out noise.
Overall, by introducing randomness in the number of correct answers, the number of supporting documents per answer, and adding misinformation and noisy documents for each query in RamDocs, we design a challenging setting simulating the variability and unpredictability of real-world retrieval scenarios. The resulting dataset comprises 500 queries in total, which on an average have 2.20 valid answers per query, and corresponding to each query, RamDocs comprises an average of 5.53 documents in the retrieved pool, out of which an average of 3.84 documents support a valid answer and the remaining 1.70 documents correspond to misinformation and noise. Empirically, we demonstrate that LLMs struggle in this challenging setting in Section 5.2 and expand on how performance degrades with each of these factors in Section 6. We refer readers to Appendix A for detailed dataset statistics and distribution plots.
@@ 4 Madam-RAG: Multi-agent Debate for Ambiguity and Misinformation in RAG
The aggregator $\mathcal{A}$ receives the outputs from all agents in round $t$, and generates a summarized answer with explanation $(y^{(t)},e^{(t)})=\mathcal{A}(\mathcal{R}^{(t)})$. The aggregator considers different agents’ responses, resolves inconsistencies, and synthesizes reliable answers based on available evidence, taking a global view over the entire conversation which allows it to take into account the source of a conflict and whether it is valid. In cases of ambiguous queries such as the example in Figure 2, where two answers refer to different individuals named “Michael Jordan”—the aggregator recognizes that both 1963 and 1956 are plausible birth years corresponding to distinct entities (the basketball player and the professor). In contrast, it flags 1998 (from the misinformation document) as factually inconsistent and irrelevant.
@@ 5.1 Experimental Setup
We evaluate Madam-RAG on three datasets: FaithEval’s inconsistent subset (Ming et al., 2024), AmbigDocs (Lee et al., 2024), and our proposed RamDocs. FaithEval tests whether LLM-generated responses remain faithful to retrieved evidence, particularly in the presence of subtle factual perturbations (i.e., misinformation) over 1000 instances. On the other hand, AmbigDocs focuses on evaluating ambiguity in multi-document settings by pairing each question with multiple documents that may support different valid answers. For AmbigDocs, we sample 1000 instances from the test set. Our RamDocs dataset is constructed to simulate real-world retrieval conditions, including multiple valid answers with varied document support, inter-document conflicts, misinformation, and noise over 500 unique queries, obtained from AmbigDocs. These datasets together enable a comprehensive assessment of model robustness under ambiguous, conflicting, and potentially misleading retrieval conditions.
@@
We test the effectiveness of Madam-RAG on two open-source post-trained language models: Llama3.3-70B-Instruct (Llama Team, 2024) and Qwen2.5-72B-Instruct (Qwen et al., 2025), as well as GPT-4o-mini (Hurst et al., 2024), representing a closed-source frontier model.
@@
For each LLM, we compare Madam-RAG with the aforementioned LLMs (for $T=3$ maximum rounds) against the following baselines (refer to prompts in Appendix C):
@@
  • Astute RAG (Wang et al., 2024a): A recent method that aligns retrieved content with the model’s parametric knowledge to resolve inconsistencies. The framework generates internal parametric knowledge document, clusters retrieved and parametric knowledge documents into consistent or conflicting groups, and then selects the most reliable cluster to generate the final answer. See details in Appendix B.
@@ 5.2 Main Results
As shown in Table 1, Madam-RAG consistently outperforms concatenated-prompt and Astute RAG baselines across all three datasets and model backbones. The gains are particularly notable on AmbigDocs and RamDocs, where ambiguity and conflicting evidence require structured resolution; for instance, Madam-RAG outperforms Astute RAG by 11.40% with Llama3.3-70B-Inst and by 12.90% with Qwen2.5-72B-Inst. On FaithEval, a dataset with conflicting evidence but only a single correct answer, Madam-RAG yields the highest performance, surpassing concatenated-prompt baseline by 15.80%, and 19.20% with Llama3.3-70B, and Qwen2.5-72B models, respectively, demonstrating its robustness to subtle misinformation. Notably, in Table 1, we find that solely using the parametric knowledge of the LLM (i.e., the No RAG baseline which does not see any of the retrieved documents) is somewhat effective on FaithEval owing to the fact that it does not have access to the misinformation in documents. However, relying on parametric knowledge alone is not sufficient as we find that LLMs when presented with ambiguous queries have the propensity to only present one valid answer, thereby, leading to poor performance on AmbigDocs and RamDocs. On the other hand, we demonstrate that Madam-RAG has the ability to ignore misinformation in retrieved documents while effectively utilizing evidence supporting multiple valid answers.
@@ 6.1 Importance of Using the Aggregator and Multiple Rounds of Debate
As shown in Figure 3, both increasing the number of debate rounds and enabling the aggregator contribute to substantial gains. Without the aggregator, performance improves substantially with more rounds (e.g., +21.10 gain in accuracy from round 1 to 3 on FaithEval and +3.62 increase in F1 on RamDocs), indicating that iterating allows agents to refine their answers and reduce error. The addition of the aggregator provides a stronger boost, especially in earlier rounds: in round 1, accuracy improves from 30.00 to 37.80 and F1 from 59.79 to 68.63 on RamDocs, reflecting the aggregator’s ability to synthesize evidence and suppress misinformation effectively. Across rounds, the aggregator yields higher precision (e.g., 67.41 $\rightarrow$ 68.52 $\rightarrow$ 68.57 for RamDocs), while recall tends to decrease slightly, suggesting a more selective and cautious answering strategy. Importantly, we argue that precision is more critical in conflict-heavy settings like RamDocs as false positives (i.e., including misinformation in the output) could be more harmful than false negatives (such as missing a correct answer). In such cases, it may be preferable for a model to be cautious and omit uncertain answers rather than risking introducing incorrect ones. The aggregator helps enforce this precision-reliability trade-off by aligning agent views early and mitigating the influence of conflicting or noisy sources. Overall, these results support the aggregator’s role in improving answer faithfulness and trustworthiness, especially under realistic, imperfect retrieval conditions.
@@ 6.3 Impact of Increasing Misinformation
Results in Figure 4(b) demonstrate that increasing the level of misinformation in retrieved evidence negatively impacts overall performance. For instance, both the concatenated-prompt and Astute RAG baselines suffer a decrease of 46%, each. While Madam-RAG also drops in performance, given a fixed degree of misinformation, Madam-RAG yields the highest performance. We note that the decrease in performance as the level of misinformation increases is to be expected, as higher degree of conflicting evidence can build mistrust around the query, making it harder for the LLM to distinguish factual documents from inaccurate ones. These findings align with those of Xu et al. (2024a), who find that even strong LLMs are susceptible to being persuaded of misinformation.

### v2

@@ 2 Related Work
Retrieval-augmented generation (RAG) has emerged as a crucial technique for enhancing language models via integrating external knowledge retrieval instead of solely relying on parametric knowledge (Lewis et al., 2020; Karpukhin et al., 2020; Cheng et al., 2021). Variants such as REALM (Guu et al., 2020), Fusion-in-Decoder (FiD) (Izacard & Grave, 2021), and RETRO (Borgeaud et al., 2022) have further optimized retrieval methods and document encoding strategies to enhance efficiency and scalability. However, these do not fully address the critical challenge of handling conflicting information across multiple retrieved documents. Prior work, such as Chen et al. (2022); Zou et al. (2024), has shown that RAG systems often propagate misinformation if retrieved documents contain errors, or arbitrarily choose an answer or use parametric knowledge to break ties when documents provide conflicting claims. Unlike existing solutions like Self-RAG (Asai et al., 2024) that use LLM-generated critiques, Astute RAG (Wang et al., 2024a) that uses the LLM’s parametric knowledge and clustering to combine retrieved documents and filter outliers or misinformation, and Speculative RAG (Wang et al., 2025), which uses a small specialist LM to draft multiple answers in parallel from distinct subset documents, and a larger generalist LM to verify and select the best one, Madam-RAG employs multi-agent debate over several rounds, wherein each agent gets the opportunity to revise its response as well as influence other agents in each round. Moreover, in contrast to Chang et al. (2024), who use a multi-agent framework solely for scoring and filtering out noisy documents, our multi-agent approach also handles other scenarios with conflicting evidence by suppressing misinformation and presenting multiple valid answers for ambiguous queries.
@@
Several benchmarks have been proposed to evaluate RAG systems under various conditions. AmbigNQ (Min et al., 2020) and AmbigDocs (Lee et al., 2024) assess the ability to handle questions with multiple valid answers, with the latter extending information about ambiguous entities across multiple documents, leading us to test on AmbigDocs. However, these datasets do not consider noise or misinformation. On the other hand, the RGB dataset (Chen et al., 2024) introduces retrieval noise, testing how well models respond to partially relevant content in the retrieved documents. HAGRID (Kamalloo et al., 2023) focuses on attribution in generative QA, while CRAG (Yang et al., 2024) presents a comprehensive benchmark covering diverse question types across five domains, varying in entity popularity and temporal sensitivity. The DRUID dataset (Hagström et al., 2025) contains internet-retrieved evidence annotated for stance and relevance, including both misinformation and insufficient or irrelevant evidence. Note that each of these datasets assumes that a given question has a single correct answer and thus they do not address scenarios with multiple conflicting yet valid answers, which our work explicitly targets. We unify these separate lines of work by introducing RamDocs that includes multiple valid answers for each query (due to ambiguity) as well as uneven numbers of supporting documents and conflicting evidence from misinformed and noisy documents.
@@
Recent studies have highlighted the challenge of knowledge conflict in LLMs, where inconsistent or conflicting information may arise from either internal parametric memory or retrieved external context. A comprehensive survey by Xu et al. (2024b) categorizes knowledge conflicts into intra-context, inter-context, and parametric conflicts, and outlines the limitations of current methods in resolving them. While prior work has made strides in addressing parametric and single-document conflicts (Gao et al., 2023; Wang et al., 2024b; Huang et al., 2025), recent studies highlight the difficulty of resolving disagreements across multiple contexts or retrieved documents (Chen et al., 2022; Su et al., 2024). Such inter-context conflict can stem from misinformation (Pan et al., 2023a; Zhou et al., 2023) or outdated knowledge (Kasai et al., 2023), and has been shown to impair factual accuracy (Jin et al., 2024). Our Madam-RAG approach addresses this gap by modeling each document by a separate LLM agent, using multi-agent debate to identify potential ambiguity and aggregating multiple valid answers while also suppressing and resolving conflicting evidence stemming from misinformation and noisy retrieval.
@@ 3 RamDocs: Retrieval with Ambiguity & Misinformation in Documents
Measuring the performance of RAG systems in real-world settings requires assessing each system’s ability to cope with inter-context conflict—disagreements among retrieved documents—which may arise due to ambiguity, misinformation, or noise. These conflict sources differ in nature: ambiguity involves different but valid answers from underspecified queries (Lee et al., 2024), misinformation refers to factually incorrect content (Ming et al., 2024), and noise includes irrelevant or weakly related documents. Moreover, different kinds of conflict demand different solutions: ambiguity should not be treated the same way as misinformation or noise; these solutions may at times conflict with each other. While existing datasets address these factors separately, in practice they will appear together and may interact in unexpected ways. This makes measuring multiple kinds of conflict in a single dataset crucial, and motivates the need for such a representative set of retrieved documents.
@@
To rigorously evaluate the robustness of current RAG systems, we present RamDocs, a dataset designed to simulate real-world retrieval challenges, including conflicting information, noise, and misinformation. Building upon AmbigDocs (Lee et al., 2024) that focuses on evaluating ambiguity by presenting multiple valid answers per query, we extend it by introducing additional real-world retrieval complexities. Specifically, we randomly sample between 1 to 3 correct answers for each ambiguous query from the dataset, resulting in 500 queries. This setup ensures that models must not only recognize ambiguity but also operate under realistic constraints where only partial information may be retrieved.
@@
Finally, for each query, we include documents containing misinformation and noise to evaluate an RAG system’s resilience against unreliable content. To incorporate misinformation, we include an additional retrieved document and replace the correct answer with a misleading or incorrect entity, ensuring the misinformation blends naturally into the document context. This is similar to the approach in Longpre et al. (2021), who replace the original answer in the context with a similar type of answer, but here, we adapt the procedure to modify the retrieved documents. We randomly determine the number of misinformation documents added to each query from 0-2, reflecting the unpredictable nature of misinformation in real-world retrieval. Additionally, we introduce noisy documents for each query (sampled from [0, 2]), which are either irrelevant to the query or contain low-quality content, to assess the system’s ability to filter out noise.
Overall, by introducing randomness in the number of correct answers, the number of supporting documents per answer, and adding misinformation and noisy documents for each query in RamDocs, we design a challenging setting simulating the variability and unpredictability of real-world retrieval scenarios. The resulting dataset comprises 500 queries, with an average of 2.20 valid answers per query. For each query, RamDocs comprises an average of 5.53 documents in the retrieved pool, out of which an average of 3.84 documents support a valid answer and the remaining 1.70 documents correspond to misinformation and noise. Empirically, we demonstrate that LLMs struggle in this challenging setting in Section 5.2 and expand on how performance degrades with each of these factors in Section 6. We refer readers to Appendix A for detailed dataset statistics and distribution plots.
@@ 4 Madam-RAG: Multi-agent Debate for Ambiguity and Misinformation in RAG
The aggregator $\mathcal{A}$ receives the outputs from all agents in round $t$, and generates a summarized answer with explanation $(y^{(t)},e^{(t)})=\mathcal{A}(\mathcal{R}^{(t)})$. To mitigate position bias, we shuffle agent responses before passing them to the aggregator. The aggregator considers different agents’ responses, resolves inconsistencies, and synthesizes reliable answers based on available evidence, taking a global view over the entire conversation which allows it to take into account the source of a conflict and whether it is valid. In cases of ambiguous queries such as the example in Figure 2, where two answers refer to different individuals named “Michael Jordan”—the aggregator recognizes that both 1963 and 1956 are plausible birth years corresponding to distinct entities (the basketball player and the professor). In contrast, it flags 1998 (from the misinformation document) as factually inconsistent and irrelevant.
@@ 5.1 Experimental Setup
We evaluate Madam-RAG on three datasets: FaithEval’s inconsistent subset (Ming et al., 2024), AmbigDocs (Lee et al., 2024), and our proposed RamDocs. FaithEval tests whether LLM-generated responses remain faithful to retrieved evidence, particularly in the presence of subtle factual perturbations (i.e., misinformation) over 1000 instances. AmbigDocs focuses on evaluating ambiguity in multi-document settings by pairing each question with multiple documents that may support different valid answers. For AmbigDocs, we sample 1000 instances from its test set. Our RamDocs is constructed to simulate real-world retrieval conditions, including multiple valid answers with varied document support, inter-document conflicts, misinformation, and noise over 500 unique queries, obtained from AmbigDocs. These datasets together enable a comprehensive assessment of model robustness under ambiguous, conflicting, and potentially misleading retrieval conditions.
@@
We test on two open-source post-trained language models: Llama3.3-70B-Instruct (Llama Team, 2024) and Qwen2.5-72B-Instruct (Qwen et al., 2025), as well as GPT-4o-mini (Hurst et al., 2024), representing a closed-source frontier model.
@@
For each LLM, we compare Madam-RAG with the aforementioned LLMs (for $T=3$ maximum rounds) against the following baselines (refer to prompts in Appendix E):
@@
  • Single Agent with Self-reflection: A three-step prompting strategy adapted from Madaan et al. (2023); Huang et al. (2024): 1) prompt LLM to generate an initial answer; 2) prompt LLM to review its previous answer and produce feedback; 3) prompt LLM to answer the question again with the feedback. We conduct two rounds of self-reflection.
  • Self-RAG (Asai et al., 2024): Self-RAG instruction-tunes an LLM to generate self-reflection tags that guide dynamic retrieval and critique the relevance of retrieved documents before answering. It processes multiple documents in parallel, producing separate responses for each. Since our setting allows multiple correct answers, we forward all responses with critique scores to the aggregator to generate the final answer, rather than selecting the highest-scoring one as in the original method.
  • Speculative RAG (Wang et al., 2025): A framework that uses a smaller, distilled specialist LM to generate drafts in parallel that are then fed to a larger generalist LM to verify and select the best draft. Each draft is generated from a distinct subset of retrieved documents, offering diverse perspectives on the evidence. See details in Appendix B.
  • Astute RAG (Wang et al., 2024a): A recent method that aligns retrieved content with the model’s parametric knowledge to resolve inconsistencies. The framework generates internal parametric knowledge document, clusters retrieved and parametric knowledge documents into consistent or conflicting groups, and then selects the most reliable cluster to generate the final answer. See details in Appendix C.
@@ 5.2 Main Results
As shown in Table 1, Madam-RAG consistently outperforms concatenated-prompt and Astute RAG baselines across all three datasets and model backbones. The gains are particularly notable on AmbigDocs and RamDocs, where ambiguity and conflicting evidence require structured resolution; for instance, Madam-RAG outperforms Astute RAG by 11.40% with Llama3.3-70B-Inst and by 12.90% with Qwen2.5-72B-Inst. On FaithEval, a dataset with conflicting evidence but only a single correct answer, Madam-RAG yields the highest performance, surpassing concatenated-prompt baseline by 15.80%, and 19.20% with Llama3.3-70B, and Qwen2.5-72B models, respectively, demonstrating its robustness to subtle misinformation. Compared to single-agent baseline with self-reflection baseline, Madam-RAG achieves gains of 17.40% on FaithEval, 3.70% on AmbigDocs, and 4.00% on RAMDocs with Llama3.3-70B-Inst. These improvements show some of the shortcomings of single-agent systems: while single-agent approaches like self-reflection (Madaan et al., 2023; Huang et al., 2024) can identify internal inconsistencies, they still process concatenated input and are therefore vulnerable to context length limits, frequency bias, and order effects (Liu et al., 2024; Zhao et al., 2021). Notably, in Table 1, we find that solely using the parametric knowledge of the LLM (i.e., the No RAG baseline which does not see any of the retrieved documents) is somewhat effective on FaithEval owing to the fact that it does not have access to the misinformation in documents. However, relying on parametric knowledge alone is not sufficient as we find that LLMs when presented with ambiguous queries have the propensity to only present one valid answer, thereby, leading to poor performance on AmbigDocs and RamDocs. On the other hand, we demonstrate that Madam-RAG has the ability to ignore misinformation in retrieved documents while effectively utilizing evidence supporting multiple valid answers.
@@
Single Agent with Self-reflection 25.70 54.50 30.40
Speculative RAG 41.80 44.30 30.60
@@
Single Agent with Self-reflection 39.40 29.40 21.80
Speculative RAG 56.20 13.40 22.20
@@
Single Agent with Self-reflection 36.00 44.50 22.50
Speculative RAG 36.50 22.50 23.00
@@ 6.1 Importance of Using the Aggregator and Multiple Rounds of Debate
As shown in Figure 3, both increasing the number of debate rounds and enabling the aggregator contribute to substantial gains. Without the aggregator, performance improves substantially with more rounds (e.g., +21.10 gain in accuracy from round 1 to 3 on FaithEval and +3.62 increase in F1 on RamDocs), indicating that iterating allows agents to refine their answers and reduce error. Adding the aggregator provides a stronger boost, especially in earlier rounds: in round 1, accuracy improves from 30.00 to 37.80 and F1 from 59.79 to 68.63 on RamDocs, reflecting the aggregator’s ability to synthesize evidence and suppress misinformation effectively. Across rounds, the aggregator yields higher precision (e.g., 67.41 $\rightarrow$ 68.52 $\rightarrow$ 68.57 for RamDocs), while recall tends to decrease slightly, suggesting a more selective and cautious answering strategy. Importantly, we argue that precision is more critical in conflict-heavy settings like RamDocs as false positives (i.e., including misinformation) could be more harmful than false negatives (e.g., missing a correct answer). In such cases, it may be preferable for a model to be cautious and omit uncertain answers rather than risking introducing incorrect ones. The aggregator helps enforce this precision-reliability trade-off by aligning agent views early and mitigating the influence of conflicting or noisy sources. Overall, these results support the aggregator’s role in improving answer faithfulness and trustworthiness, especially under realistic, imperfect retrieval conditions.
@@ 6.3 Impact of Increasing Misinformation
Results in Figure 4(b) show that increasing the level of misinformation in retrieved evidence negatively impacts overall performance. For instance, both concatenated-prompt and Astute RAG baselines suffer a decrease of 46%, each. While Madam-RAG also drops in performance, given a fixed degree of misinformation, Madam-RAG yields the highest performance. This robustness stems in part from how evidence is processed: processing all documents jointly may obscure low-frequency but valid evidence, especially when misinformation or noise dominates. Our multi-agent setup encourages each document to be independently summarized and defended, allowing underrepresented but correct views to be preserved and strengthened through debate. We note that the performance degradation with more misinformation is to be expected, as higher degree of conflicting evidence can build mistrust around the query, making it harder for the LLM to distinguish factual documents from inaccurate ones. These findings align with those of Xu et al. (2024a), who find that even strong LLMs are susceptible to being persuaded of misinformation.